OpenStudy (sphott51):

independent or dependent and probability.. the probability that two cards drawn from a deck are both aces

OpenStudy (perl):

what is the probability that the first card is an ace?

OpenStudy (sphott51):

1/52?

Directrix (directrix):

How many aces are in the deck? @SPHott51

OpenStudy (sphott51):

4

Directrix (directrix):

The probability of getting the first Ace is 4/52 because there are 4 ways to get an Ace and there are 52 cards in the deck.

Directrix (directrix):

When you draw the second card to see if you can get an Ace to go with the first one Ace you got, there are now only 3 Aces left in the deck. How many cards are in the deck when you draw for the second Ace?

Directrix (directrix):

@SPHott51

OpenStudy (sphott51):

3?

Directrix (directrix):

There are 3 Aces in the deck ready for the second draw and 51 cards in the deck. >>independent or dependent and probability. The events are dependent because the probability of drawing a second Ace is altered by the drawing of the first Ace. P(A and A) = 4/52 * 3/51 = ?

Directrix (directrix):

Dependent P(A and A) = 4/52 * 3/51 = ? 4/52 times 3/51 is the probability. Do that calculation and you'll be finished with this. @SPHott51